2x+5=0 solve it answer isnot -5/2 if you show the same answer you will get dislike
Answer #1


Given that equation 81+520 Let us taken gu+5zo overzz 20+2=0 [5=2 overz, an=-2 n=1 overz] Elza d=2 [+1=2-1=2 CS 7=2 fs. solut